Hi there could anyone help me with questions and answer it on Excel with full explantion please.

You've taken an internship with the Moscow Police Department and are analyzing crime rates to see if MPD wants to shift
patrols from a low crime area to an increasing crime area.
Month Crimes Sample
Size
January 7 1000
February 9 1000
March 7 1000
April 7 1000
May 7 1000
June 9 1000
July 7 1000
August 10 1000
September 8 1000
October 11 1000
November 10 1000
December 8 1000
a. Construct a p chart using a 95% confidence level (2 standard deviations above/below the mean) and plot the year's
data
b. You've collected data for the first three months of the following year. In analyzing this data, what comments can you
make regarding the crime rate?
Crime Incidence Sample Size
January 10 500
February 12 500
Marcy 11 500
Answer #1

For a p chart

pbar = no of defects / no of samples x sample size =100 /12x1000 = 1/120 = 0.00833

UCL = npbar + 2 [ npbar ( 1-pbar)/n]1/2 = 8.33 + 2 [ 8.33 ( 1-0.00833)/1000]1/2 = 8.33+0.18 = 8.51

LCL =  npbar - 2 [ npbar ( 1-pbar)/n]1/2 = 8.33 -0.18 = 8.15

Chart

to beb. From the first three months' observation, it can be concluded that the crime rate has increased manifold, with 20, 14 and 22 in the firs three months of the year in comparison to what was 7,9 and 7 last year during these months.